OpenStudy (ximstorm):

Which expression is equal to 4y? y + 3 y2 + y2 3y + 1 y + y + y + y

jimthompson5910 (jim_thompson5910):

y + y + y + y = 4y you can think of it as 1y + 1y + 1y + 1y = (1+1+1+1)y = 4y or as (1 yak) + (1 yak) + (1 yak) + (1 yak) = 4 yaks

jimthompson5910 (jim_thompson5910):

y + 3 would NOT turn into 4y because the two terms (y and 3) are NOT like terms you can't say y+3 = 4y or you can't say y+3 = 4. You just leave it as y+3 and you can't simplify any further. The same applies to the other expressions except for the last one